OpenStudy (anonymous):

Solve by using elimination -2x+2y+3z=0 -2x-y+z=-3 2x+3y+3z=5 Solve by using substitution -x-y-z=-8 -4x+4y+5z=7 2x+2z=4

rishavraj (rishavraj):

for elimination on just subtract the first two equations..... x gets eliminated ...... then proceed

rishavraj (rishavraj):

@Rwelch ??

OpenStudy (anonymous):

well i tried that but i had gotten the answer wrong but teacher didnt tell me how to fix it

rishavraj (rishavraj):

after doing the manipulation i told u will get an equation in y and z 3y + 2z = 3 then add second and third equation...x gets weliminated..and u get another equation in y and z

OpenStudy (anonymous):

um what ?

rishavraj (rishavraj):

see the first equation is -2x+2y+3z=0 second equation is -2x-y+z=-3 third equation is 2x+3y+3z=5 now simply add equation 1 and 3 x gets eliminated and again add equation 1 and 3 x gets eliminated... u will hav two equation in y and z then get values of y and z ang plug them in any of the given three equation

OpenStudy (anonymous):

nope still not understanding

rishavraj (rishavraj):

oops my bad